OpenStudy (anonymous):

What Does X Equals in 95 (x+2) = 145 ?

OpenStudy (anonymous):

x=-9/19

hero (hero):

@Siddharth18 , at least provide some kind of explanation BEFORE giving the answer.

hero (hero):

If the person is already present, you could preferably try LEADING him or her to the answer.

OpenStudy (anonymous):

x+ 2=145/95=29/19 x=29/19-2= -9/19